OpenStudy (anonymous):

The sum of a number and teice its reciprocal is 3. Find the number

OpenStudy (whpalmer4):

let n be the number \[3 = n + 2(\frac{1}{n})\] Multiply both sides by n\[3n=n^2+2\]subtract 3n from both sides\[n^2-3n+2=0\]Factor or use the quadratic formula to solve for n. Hint: there may be multiple numbers that work!

OpenStudy (anonymous):

I got 1 and 2 is that right?

OpenStudy (whpalmer4):

Well, let's check: 1+2*(1/1) = 1+2 = 3 check 2+2*(1/2) = 2+1 = 3 check Because it was a quadratic, it has two solutions, and you found them!